OpenStudy (anonymous):

What is the slope intercept form that passes through the given point and is perpendicular (-6,4), 3y =2x-3? What is the answer?! Help!! :)

OpenStudy (anonymous):

slope-intercept form is y=mx+b so you're going to want to take the full equation that you have and divide it by 3 3y=2x-3 becomes y=(2/3)x-1 your slope (m) is 2/3, the slope of a line perpendicular to this one is the negative reciprocal of this slope, which means you flip it and multiply it by -1 the slope of a line perpendicular to this one is then: -3/2 Now you take your x, y, and m values: x=-6 y=4 m=-3/2 and plug them into the slope-intercept form 4=(-3/2)(-6) + b multiply -3/2 by -6 to get 9 (remember, a negative times a negative is always a positive) you now have 4=9+b subtract nine from both sides to get b by itself -5=b Now re-write the equation with only m and b, leaving x and y as variables: y = (-3/2)x - 5